How many grams of sodium phosphate, Na3PO4, must be dissolved to prepare 400. mL of a 0.182 M aqueous solution of the salt?
Answer #1

Form the general equation of Molarity, we solve for moles and then, the mass:

M = moles / V

moles = m/MM

Replacing this, in the expression of molarity:

M = m / MM * V solving for mass:

m = MM * V * M

The MM is: (3*23) + 31 + 16*4 = 164 g/mol

m = 164 g/mol * 0.400 L * 0.182 mol/L

m = 11.9392 g

Hope this helps